# Examining dark neutron decay using f-mode gravitational waves

### v1

@@ 8 F-mode frequency
The frequency of an f-mode is highly sensitive to the EoS of the ultra-dense matter within the neutron star [112, 113, 114]. A stiffer EoS (less compressible) generally predicts higher f-mode frequencies. Conversely, a softer EoS (more compressible) predicts lower f-mode frequencies. The acceptable values of f-mode frequency for compact stars, such as neutron stars, typically range between 1 and 1–3 kHz. These frequencies are significant because they fall within the detection range of advanced gravitational wave detectors like the Einstein Telescope and Cosmic Explorer [23]. Figure (7) illustrates that the f-mode frequency of the selected models adheres to the universal relation, demonstrating consistency across different astrophysical scenarios. Notably, both frequency ranges fall within the detection capabilities of the Einstein Telescope and Cosmic Explorer, highlighting their potential for observational verification. Furthermore, for neutron decay into dark matter, variations in dark matter self-interaction strength do not disrupt the universal trend. Regardless of the DM-DM self-interaction parameter, the f-mode frequency remains well within the expected range, reinforcing the robustness of the underlying physical principles governing these oscillations. This consistency suggests that such models could be effectively probed through next-generation gravitational wave observatories.
@@ 9 Conclusion
In conclusion, this study highlights the impact of different equations of state (EoSs) on neutron star properties, including mass, radius, damping time, and oscillation frequencies. The findings confirm that nucleon-only EoSs produce the stiffest models, while increasing dark matter (DM) self-repulsion strengthens the equation of state. Notably, neutron decay into dark matter aligns more closely with nucleon-only EoSs as self-interaction increases.
The study further reinforces the existence of universal relations in neutron star oscillations. Both the first- and second-order damping time equations demonstrate consistency across selected EoSs, with strange matter and non-self-interacting dark matter exhibiting deviations. Similarly, mass-scaled angular frequency correlates with compactness, suggesting that more compact neutron stars may emit stronger gravitational waves, making them promising candidates for detection.
Additionally, the f-mode frequency results indicate that all selected models follow a universal trend, remaining within the detection range of next-generation observatories such as the Einstein Telescope and Cosmic Explorer. This consistency supports the viability of probing neutron star interiors through gravitational wave observations, emphasizing the importance of understanding the role of dark matter interactions in dense astrophysical objects.
The results of this study enhance our understanding of neutron star behavior and provide valuable insights for future gravitational wave detections, opening avenues for further investigations into exotic matter compositions and their astrophysical implications.

### v2

@@ 8 F-mode frequency
The frequency of an f-mode is highly sensitive to the EoS of the ultra-dense matter within the neutron star [116, 117, 118]. A stiffer EoS generally predicts higher f-mode frequencies. Conversely, a softer EoS predicts lower f-mode frequencies. The acceptable values of f-mode frequency for compact stars, such as neutron stars, typically range between 1 and 1–3 kHz. These frequencies are significant because they fall within the detection range of advanced gravitational wave detectors like the Einstein Telescope and Cosmic Explorer [24]. Figure (6) illustrates that the f-mode frequency of the selected models adheres to the universal relation, demonstrating consistency across different astrophysical scenarios. Notably, both frequency ranges fall within the detection capabilities of the Einstein Telescope and Cosmic Explorer, highlighting their potential for observational verification. Furthermore, for neutron decay into dark matter, variations in dark matter self-interaction strength do not disrupt the universal trend. Regardless of the DM-DM self-interaction parameter, the f-mode frequency remains well within the expected range, reinforcing the robustness of the underlying physical principles governing these oscillations. This consistency suggests that such models could be effectively probed through next-generation gravitational wave observatories.
Our results are consistent with earlier works exploring neutron-to-dark matter decay in neutron stars. Motta et al. [42] and Grinstein et al. [17] found that non-interacting dark fermions destabilize neutron stars, reducing their maximum mass below the threshold, which we confirm in our analysis. Husain et al. [50] similarly emphasized the necessity of DM self-repulsion, consistent with our finding that fm2 restores compatibility with observational constraints.
For f-mode oscillations, our study extends the universal relations established by Andersson and Kokkotas [125] and refined by Lioutas and Stergioulas [126], showing that these relations persist even in the presence of exotic matter. However, we demonstrate that strange matter and non-self-interacting DM produce systematic deviations, suggesting that precise future detections could be used to discriminate between EoSs.
@@ 9 Conclusion
In this work, we investigated the impact of nucleonic, strange matter, and dark matter equations of state on the f-mode oscillation properties of neutron stars, with particular attention to damping times and universal relations. Our results demonstrate that the nucleon-only EoS is consistently the stiffest, while the inclusion of dark matter generally softens the stellar structure, unless significant DM–DM self-interactions are present. Increasing the self-repulsion parameter suppresses dark matter production, thereby stiffening the EoS and bringing the neutron-decay scenario closer to the nucleonic case. In contrast, strange matter exhibits longer damping times and notable deviations at low densities, reflecting its softer pressure support relative to nucleonic matter.
Importantly, we find that neutron decay into dark matter leads to stellar configurations that remain compatible with current astrophysical constraints. Moreover, the corresponding f-mode frequencies of such stars fall squarely within the detection range of modern gravitational-wave interferometers, underscoring their observational relevance.
Despite these compositional differences, the mass-scaled angular frequency retains quasi-universal behavior when expressed as a function of compactness () and the structural parameter . This universality is preserved even in the presence of exotic matter, with only minor deviations in the case of strange matter and non-self-interacting dark matter. The persistence of such relations highlights their robustness and utility as probes of neutron star structure.
Astrophysically, these results suggest that f-mode oscillations and their associated gravitational-wave signals can provide meaningful constraints on neutron star properties without requiring precise knowledge of the microscopic EoS. At the same time, the systematic deviations observed in exotic matter scenarios offer potential observational signatures that could help discriminate between standard nucleonic models and stars containing dark or strange matter. Future detections of f-mode gravitational waves by next-generation interferometers could therefore serve as a powerful testbed for probing the presence of exotic matter in compact stars and for constraining the strength of dark matter self-interactions. In particular, estimating the signal-to-noise ratios of the predicted f-modes in next-generation detectors such as the Einstein Telescope and Cosmic Explorer, for realistic excitation energies and astrophysical channels (post-merger remnants, proto-neutron stars, or glitches), will be an important direction. Such studies could establish whether the parameter space explored here can be probed observationally, and whether stacking strategies or Bayesian inference using universal relations can enhance detectability. This will be pursued in future work.